# A Survey on Memory-Efficient Large-Scale Model Training in AI for Science

### v1

@@ Abstract
Scientific research faces high costs and inefficiencies with traditional methods, but the rise of deep learning and large language models (LLMs) offers innovative solutions. This survey reviews LLM applications across scientific fields such as biology, medicine, chemistry, and meteorology, underscoring their role in advancing research. However, the continuous expansion of model size has led to significant memory demands, hindering further development and application of LLMs for science. To address this, we review memory-efficient training techniques for LLMs based on the transformer architecture, including distributed training, mixed precision training, and gradient checkpointing. Using AlphaFold 2 as an example, we demonstrate how tailored memory optimization methods can reduce storage needs while preserving prediction accuracy. We also discuss the challenges of memory optimization in practice and potential future directions, hoping to provide valuable insights for researchers and engineers.
Index Terms:
AI for science, memory optimization, large-scale deep learning model, distributed training.
I Introduction
The rapid advancement of artificial intelligence, particularly large language models (LLMs), has positioned deep learning (DL) as a vital tool for addressing various scientific challenges. DL offers a new approach to scientific research due to its strengths in automatic feature extraction and complex pattern recognition. According to the universal approximation theorem, deep neural networks (DNNs) can effectively fit any continuous function. Additionally, the high-performance parallel implementation of the backpropagation algorithm on GPUs enables DL to efficiently solve scientific and engineering problems [1]. In disciplines such as biology, medicine, chemistry, and climate science, DL models demonstrate great potential to assist researchers in accelerating scientific discoveries, enhancing prediction accuracy, and providing innovative solutions to complex challenges.
In Biological Sciences, it has proven effective in areas such as protein structure and function prediction, as well as genetic engineering. One notable example is AlphaFold 2 [2], which uses DL to predict the three-dimensional structures of proteins. In the medical field, LLMs have been applied to various tasks, including medical Q&A, disease prediction, and clinical decision support. For instance, the Med-PaLM series of models [3], [4], [5] can answer questions posed by medical professionals and assist doctors in diagnosing diseases and determining treatment options. In chemistry, DL is employed for tasks such as molecular generation [6], [7], molecular property prediction [8], and chemical reaction prediction [9]. In meteorology, by processing vast amounts of meteorological data, DL models can reveal the complex laws of weather change, improve the accuracy of weather prediction, simulate and predict a wide range of climate patterns [10], [11], [12], and predict the occurrence and impact of extreme weather events [13], [14]. With the improvement of computing power and the continuous progress of DL algorithms, the application of DL in scientific fields will be more extensive and in-depth, further expanding the boundaries of scientific exploration.
@@ I Introduction
Meanwhile, studies have shown that with the increase of parameters, the model shows better performance [15], [16]. Consequently, the scale of these models has continued to grow, leading to memory usage issues in DL models. In recent years, the widespread development and application of transformer models [17] have led to a significant increase in model size, averaging a 240-fold growth every two years. The expansion of parameters in models like GPT-3 [18] and Llama 2 [19] has outpaced the growth of GPU memory. The training of large-scale models inevitably encounters the problem of the “memory wall” [20] (see Fig. 1).
@@
There are a few related surveys summarized memory-efficient LLM training methods. Zhuang et al. [21] have provided an overview of the efficient training technologies used for transformer-based models. They discussed some memory-efficient training methods and summarized the key factors that affect training efficiency. To gain a deeper understanding of the latest advancements in efficient LLMs, Wan et al. [22] conducted a thorough literature review and proposed a taxonomy to organize the relevant research. In contrast to these studies, this survey focuses on the application of LLMs across different scientific fields, systematically introducing general memory-efficient training methods for transformer-based large-scale scientific models. Furthermore, we use a case study to demonstrate studies aimed at optimizing memory overhead in transformer-variants-based models. To the best of our knowledge, there are no existing reviews specifically addressing memory-efficient training methods for LLMs in scientific fields.
@@
Our investigation indicates that despite the widespread application of LLMs in scientific fields, there is a notable lack of memory efficiency in their training processes. A judicious selection of memory optimization techniques is essential, as it can significantly improve the efficiency of DL models and broaden their applicability in resource-constrained settings. It is of considerable theoretical and practical importance for facilitating the broader adoption of DL technologies in scientific research.
@@
The remainder of this paper is organized as follows. In Section II, we begin by providing a brief introduction to deep learning and the associated knowledge of LLMs to enhance understanding for researchers and practitioners from various fields. We then highlight the accomplishments of LLMs across different scientific domains and summarize the memory optimization techniques employed during model training. Section III offers a comprehensive overview of various strategies aimed at reducing memory overhead during LLM training. Section IV focuses on AlphaFold 2 as a representative example, detailing some customized memory optimization methods. Section V addresses the trends and challenges of applying memory optimization techniques in large-scale models within scientific contexts. Finally, Section VI reviews and summarizes the key points discussed throughout the paper.
II Large Language Models and AI for Science
II-A A Brief Introduction to Deep Learning and LLMs
@@
LLMs are DNN-based language models characterized by a massive number of parameters. These models enhance performance by increasing both capacity and the volume of training data. Before the rise of deep learning, language modeling primarily relied on statistical methods, such as n-gram models and hidden Markov models. However, they struggled with data sparsity, making it challenging to handle long sequences and complex dependencies effectively. In 2003, Bengio et al. [23] introduced a neural network-based language model that sparked a wave of research into neural language models. Subsequently, Mikolov et al. proposed Word2Vec, which enhanced the representation of words through Skip-gram [24] and CBOW [25] methods. This innovation became a vital tool for feature representation in later natural language processing (NLP) research and applications. During this period, models like recurrent neural network (RNN) [26], long short-term memory network (LSTM) [27], and gated recurrent unit (GRU) [28] gained popularity for language modeling because they could better capture long-distance dependencies. Despite this progress, these approaches still encounter efficiency limitations when processing long sequences, and their training processes are challenging to parallelize.
@@
The proposal of Transformer architecture [17] in 2017 marked a revolutionary breakthrough in the field of language modeling. Transformers use a multi-head self-attention mechanism rather than a recurrent structure, enabling them to efficiently capture global dependencies in long sequences and enhance model training speed. With these advancements, transformers quickly became the foundation of modern NLP models, driving the rapid evolution of LLMs. Models like BERT [29] and GPT [30] have emerged one after another, establishing pre-training and fine-tuning as an NLP paradigm. Subsequently, GPT-2 [31] and GPT-3 [18] expanded their parameter sizes to 1.5B and 175B, demonstrating the potential for large-scale pre-training, particularly in zero-shot and few-shot learning. The Scaling Laws [15], [16] reveal the relationship between the growth of model parameters, data size, computational complexity, and model performance, providing theoretical guidance for designing and training large-scale models. To further improve the performance of the model on tasks, researchers have also proposed methods such as instruction fine-tuning [32] and InstructGPT [33]. At the end of 2022, ChatGPT [^1] https://openai.com/index/chatgpt showcased the tremendous potential of LLM to the world and sparked a research boom in LLM.
@@
In recent years, companies and research institutions have introduced a range of LLMs, advancing their scale and capabilities. Notable examples include but not limited to Google’s PaLM series [34], [35], Meta’s Llama series [36], [19], [37], OpenAI’s GPT-4 [38], DeepSeek-AI’s DeepSeek series [39], [40], [41], Tsinghua’s GLM-4 [42], and Alibaba’s Qwen series [43], [44]. So far, LLMs have exhibited exceptional proficiency in language generation, multimodal integration, and specialized tasks, providing substantial support for scientific research, industrial applications, and everyday life. As technological advancements continue, LLMs are expected to exert a significant influence across a wider range of domains.
II-B Application of LLMs in Scientific Fields
@@
With the success of LLMs in NLP, an increasing amount of research is exploring its application to scientific fields. As a powerful tool for knowledge expression and reasoning, LLMs are capable of managing complex high-dimensional data, uncovering deep patterns, and enhancing the speed and accuracy of scientific discoveries. This section will focus on the use of LLMs in common tasks across different scientific disciplines. By reviewing these applications, we aim to illustrate the broad applicability and impressive capabilities of LLMs in diverse areas of science, as well as the necessity of memory-efficient training techniques. Table I summarizes relevant studies utilizing LLMs across multiple fields and outlines the memory optimization techniques employed during model training. It is noteworthy that as activation memory usage varies with experiment settings, the straightforward estimation of memory cost in Table I does not include activations, which constitute a predominant portion of memory overhead during the training process of certain models.
@@
TABLE I: Applications of LLMs in Scientific Fields and Corresponding Training Optimization Strategies
@@
  Biology AlphaFold 2 [2] - Evoformer [2] 93M 1.45 GB DP mixed-precision GC protein structure prediction
RosettaFold [45] - SE(3)-Transformer [46] 130M 2.03 GB DP GA protein structure prediction
AlphaFold 3 [47] - Pairformer [47] Unreported - Unreported biomolecular complex structure prediction
OpenFold [48] - Evoformer 93M 1.45 GB DP (ZeRO-2) mixed-precision GC offloading GA protein structure prediction
FastFold [49] - Evoformer 93M 1.45 GB DP DAP protein structure prediction
ScaleFold [50] - Evoformer 97M 1.52 GB DP DAP protein structure prediction
ESMFold [51] ESM-2 15B Transformer [17] 15B 240 GB DP (FSDP) protein structure prediction
xTrimoPGLM [52] xTrimoPGLM-100B Transformer 100B 1.56 TB DP (ZeRO-1) PP (1F1B) TP (Megatron-LM) mixed-precision GC protein understanding protein generation
ESM3 [53] - Transformer 1.4B / 7B / 98B 1.53 TB DP (FSDP) mixed-precision protein reasoning protein generation
Medicine BioGPT [54] GPT-2 Medium [31] Transformer 347M 5.42 GB DP GA biomedical text understanding biomedical text generation
Med-PaLM [3] PaLM 540B [34] Flan-PaLM 540B [32] Transformer 540B 8.44 TB DP (ZeRO-3) TP GC medical question answering medical reasoning
Med-PaLM 2 [4] PaLM 2 340B [35] Transformer 340B 5.31 TB Unreported medical question answering medical reasoning
Med-PaLM M [5] PaLM-E [55] Transformer 12B / 84B / 562B 8.78 TB DP (ZeRO-3) TP GC biomedical generalist
BiomedGPT [56] OFA [57] Transformer 33M / 93M / 182M 2.84 GB DP (PyTorch DDP) mixed-precision biomedical generalist
Meditron [58] Llama-2 [19] Transformer 7B / 70B 1.09 TB DP PP TP (Megatron-LM) medical reasoning
HuatuoGPT [59] Baichuan-7B Ziya-LLaMA-13B [60] Transformer 7B / 13B 208 GB DP (ZeRO) medical consultation
HuatuoGPT-II [61] Baichuan2 [62] Yi-34B [63] Transformer 7B / 13B / 34B 544 GB DP (ZeRO) medical consultation
@@
PharmGPT [65] Llama-2 Transformer 3B / 13B / 70B 1.09 TB DP+PP+TP text understanding text generation
Chemistry ChemBERT [66] BERT-Base Transformer 110M 1.72 GB Unreported product extraction reaction role labeling
@@
Chemformer [9] BART [68] Transformer 45M / 230M 3.59 GB DP (ZeRO-2) mixed-precision reaction prediction molecular optimization molecular property prediction
MolGen [7] BART Transformer 355M 5.55 GB DP (ZeRO-2) mixed-precision GA molecule generation
@@
FuXi [12] - Swin transformer [71] 4.5B 72 GB DP (FSDP) mixed-precision GC weather forecast
ClimaX [13] - ViT [72] Unreported - DP mixed-precision weather forecast climate projection
Aurora [14] - Swin transformer 1.3B 20.80 GB DP mixed-precision GC atmospheric prediction
  • GA: gradient accumulation
  • GC: gradient checkpointing
II-B1 Biology
@@
Protein structure prediction stands as one of the most significant challenges in biochemistry, as high-precision predictions are essential for drug discovery and protein design. AlphaFold 2 [2], developed by DeepMind, is a DL-based tool for predicting protein structures. It can predict the three-dimensional structure of proteins based on their amino acid sequences, representing a major breakthrough in addressing long-standing biological challenges.
@@
The architecture of AlphaFold 2 is built upon variants of Transformer. Its main structure features an input module that receives the amino acid sequence of the target protein and searches for homologs within a sequence database to perform multiple sequence alignment (MSA). The Evoformer module then processes the MSA representation and pair representation, facilitating the exchange and updating of information through Transformer layers. Finally, the structural module converts the abstract representation of proteins into 3D atomic coordinates [73]. AlphaFold 2 has significantly influenced the advancement of biotechnology. However, its training process and datasets are not fully open source, which restricts researchers from applying it to tasks beyond protein structure prediction. Furthermore, despite the AlphaFold 2 model having a relatively small parameter size (about 93M), the memory consumed by intermediate activations is substantial due to its architecture, and peak memory usage increases cubically with the length of the input sequence. Despite using mixed precision and gradient checkpointing, the memory cost of activations in attention module exceeds 20 GB [49]. To facilitate further research, Ahdritz et al. developed OpenFold [48]. As an open-source implementation of AlphaFold 2, it achieves comparable prediction accuracy while runing faster and consuming less memory. It is built based on PyTorch [74], which makes it easier for the community to use and improve.
@@
Although AlphaFold2 and RoseTTAFold [45] achieved breakthroughs in atomic resolution structure prediction, they rely on MSA and similar protein structure templates to achieve better performance. Based on the research of OpenFold, Lin et al. [51] further explored the ability of language model and proposed ESMFold. ESMFold utilizes the internal representation of language model to replace the expensive network module for processing MSA with a transformer module for processing sequences. This simplification greatly improves the speed of ESMFold, which is much higher than models based on MSA. To enable the training of ESMFold, ZeRO stage 3 [75] is used to divide the model parameters, gradients, and optimizer states among GPUs.
@@
The protein language model (PLM) has achieved remarkable success in learning biological information from protein sequences. Most existing models are constrained by autoencoding or autoregressive pre-training objectives, making it challenging to address both protein understanding and generation tasks. Chen et al. [52] explored the possibility of jointly optimizing these two training objectives through an innovative framework and trained a 100B PLM (xTrimoPGLM). Inspired by ESMFold, they combined the folding module with xTrimoPGLM and developed xT-Fold, a high-performance 3D structure prediction tool. The performance of xT-Fold is superior to other PLM-based models on benchmarks such as CAMEO and CASP15. As the model size and the scale of training data are greatly increased, xTrimoPGLM uses various memory optimization methods in the pre-training stage, including 3D-parallelism [76], [77], [75], FP16 mixed precision training [78], and gradient checkpointing [79].
@@
In May 2024, Isomorphic Labs and DeepMind jointly released AlphaFold 3 [47]. The overall architecture of Alphafold 3 is similar to AlphaFold 2, its prediction accuracy exceeds AlphaFold 2, and the model is more efficient. In addition, the application scope of Alphafold 3 has also been expanded. It can predict the structure of a single protein, the complex structure of protein complexes, proteins, nucleic acids, small molecules, etc.
II-B2 Medicine

### v2

# A survey on memory-efficient transformer-based model training in AI for science

@@ Abstract
Scientific research faces high costs and inefficiencies with traditional methods, but the rise of deep learning and large language models (LLMs) offers innovative solutions. This survey reviews transformer-based LLM applications across scientific fields such as biology, medicine, chemistry, and meteorology, underscoring their role in advancing research. However, the continuous expansion of model size has led to significant memory demands, hindering further development and application of LLMs for science. This survey systematically reviews and categorizes memory-efficient pre-training techniques for large-scale transformers, including algorithm-level, system-level, and hardware-software co-optimization. Using AlphaFold 2 as an example, we demonstrate how tailored memory optimization methods can reduce storage needs while preserving prediction accuracy. By bridging model efficiency and scientific application needs, we hope to provide insights for scalable and cost-effective LLM training in AI for science.
keywords:
AI for Science, Memory Optimization, Large Language Model, Distributed Training.
[…]
1 Introduction
The rapid advancement of artificial intelligence, particularly large language models (LLMs), has positioned deep learning (DL) as a vital tool for addressing various scientific challenges. DL offers a new approach to scientific research due to its strengths in automatic feature extraction and complex pattern recognition. According to the universal approximation theorem, deep neural networks (DNNs) can effectively fit any continuous function. Additionally, the high-performance parallel implementation of the backpropagation algorithm on GPUs enables DL to solve scientific and engineering problems efficiently[1]. DL models demonstrate great potential to assist researchers in accelerating scientific discoveries, enhancing prediction accuracy, and providing innovative solutions to complex challenges.
@@ 1 Introduction
In Biological Sciences, DL approaches have proven effective in areas such as protein structure and function prediction. One notable example is AlphaFold 2 [2], which predicts the three-dimensional structures of proteins. In the medical field, LLMs have been applied to various tasks, including medical Q&A, disease prediction, and clinical decision support. For instance, the Med-PaLM series of models [3, 4, 5] can answer questions posed by medical professionals and assist doctors in diagnosing diseases and determining treatment options. In chemistry, DL is employed for tasks such as molecular generation [6, 7], molecular property prediction [8], and chemical reaction prediction [9]. In meteorology, by processing vast amounts of meteorological data, DL models can reveal the complex laws of weather change, improve the accuracy of weather prediction, simulate and predict a wide range of climate patterns [10, 11, 12], and predict the occurrence and impact of extreme weather events [13, 14]. With the improvement of computing power and DL algorithms, the application of DL in scientific fields will be more extensive, further expanding the boundaries of scientific exploration.
@@
Meanwhile, studies have shown that with parameters increase, the model shows better performance [15, 16]. Consequently, the scale of these models has continued to grow, leading to memory usage issues in DL models. In recent years, the widespread development and application of transformer models [17] have led to a significant increase in model size, averaging a 240-fold growth every two years. The expansion of parameters in models like GPT-3 [18] and Llama 2 [19] has outpaced the growth of GPU memory. The training of large-scale models inevitably encounters the “memory wall” [20] (see Fig. 1). Addressing this challenge is critical for enabling the training of large-scale models in resource-constrained environments.
@@
There are a few related surveys summarized memory-efficient LLM training methods. Zhuang et al. [21] have provided an overview of the efficient training technologies used for transformer-based models. They discussed some memory-efficient training methods and summarized the key factors that affect training efficiency. To gain a deeper understanding of the latest advancements in efficient LLMs, Wan et al. [22] conducted a thorough literature review and proposed a taxonomy to organize the relevant research.
In contrast to these studies, this survey investigates transformer-based models applied across different scientific fields and their corresponding memory-efficient pre-training techniques. Moreover, we classify them into existing base models, and systematically introduce and categorize trending general memory-efficient pre-training methods, thereby bridging the gap between memory-efficient training methodologies and the need for scaling scientific models. Furthermore, we use a case study to demonstrate studies aimed at optimizing memory overhead in transformer-variants-based models. To the best of our knowledge, there is a notable absence of a comprehensive review that summarizes the application of memory-efficient training techniques within the domain of AI for Science, this survey is the first to systematically connect memory-efficient training techniques to LLMs in AI for science.
@@
Our investigation indicates that despite the widespread application of transformers in scientific fields, there is a notable lack of memory efficiency in their training processes. A judicious selection of memory optimization techniques is essential, as it can significantly improve the efficiency of DL models and broaden their applicability in resource-constrained settings. It is of considerable theoretical and practical importance for facilitating the broader adoption of DL in scientific research.
@@
The remainder of this paper is organized as follows. In Section 2, we begin by providing a brief introduction to deep learning and the associated knowledge of LLMs to enhance understanding for researchers and practitioners from various fields. We then highlight the accomplishments of LLMs across different scientific domains and summarize the memory optimization techniques employed during model pre-training. Section 3 offers a comprehensive overview of various strategies aimed at reducing memory overhead. Section 4 focuses on AlphaFold 2 as a representative example, detailing some customized memory optimization methods. Section 5 addresses the trends and challenges of applying memory optimization techniques in large-scale models within scientific contexts. Finally, Section 6 reviews and summarizes the key points discussed throughout the paper.
2 Large Language Models and AI for Science
2.1 A Brief Introduction to Deep Learning and LLMs
@@
The proposal of the Transformer architecture [17] marked a revolutionary breakthrough in the field of language modeling. Transformers use a multi-head self-attention mechanism, enabling them to efficiently capture global dependencies in long sequences and enhance model training speed. Models like BERT [23] and GPT [24] have emerged one after another, establishing pre-training and fine-tuning as a standard paradigm in natural language processing (NLP). Subsequently, GPT-2 [25] and GPT-3 [18] expanded their parameter sizes to 1.5B and 175B, demonstrating the potential for large-scale pre-training, particularly in zero-shot and few-shot learning. The Scaling Laws [15, 16] reveal the relationship between the growth of model parameters, data size, computational complexity, and model performance, providing theoretical guidance for designing and training large-scale models. To further improve the performance of the model on tasks, researchers have also proposed methods such as instruction fine-tuning [26] and InstructGPT [27]. At the end of 2022, ChatGPT showcased the tremendous potential of LLM to the world and sparked a research boom in LLM.
@@
In recent years, companies and research institutions have introduced a range of LLMs, advancing their scale and capabilities. Notable examples include but not limited to Google’s PaLM series [28, 29], Meta’s Llama series [30, 19, 31], OpenAI’s GPT-4 [32], DeepSeek-AI’s DeepSeek series [33, 34, 35], Tsinghua’s GLM-4 [36], and Alibaba’s Qwen series [37, 38]. So far, LLMs have exhibited exceptional proficiency in language generation, multimodal integration, and specialized tasks, providing substantial support for scientific research, industrial applications, and everyday life.
2.2 Application of LLMs in Scientific Fields
@@
With the success of LLMs in NLP, an increasing amount of research is exploring its application to scientific fields. This section will focus on the use of transformer-based LLMs in common tasks across different scientific disciplines as summarized in Table 1. Across domains, medical models (e.g., Med-PaLM [3], Meditron [39]) emphasize instruction tuning and expert QA performance; chemistry-oriented models (e.g., Chemformer [9], MolGen [7]) focus on molecule generation and reaction prediction; climate and Earth science models deal with long-sequence and spatiotemporal data modeling, often requiring architecture adaptation. These models vary in size, training data, and optimization techniques, reflecting distinct memory challenges in each scientific domain task. By reviewing these applications, we aim to illustrate the broad applicability and capabilities of LLMs, as well as the necessity of memory-efficient training techniques. Table 1 outlines the memory optimization techniques employed during model training. It is noteworthy that as activation memory usage varies with experiment settings, the straightforward estimation of memory cost in Table 1 does not include activations, which constitute a predominant portion of memory overhead during the training process of certain models.
@@
Table 1: Applications of LLMs in scientific fields and corresponding training optimization strategies
@@
Biology AlphaFold 2 [2] - Evoformer [2] 93M 1.45 GB
DP
mixed-precision
GC
protein structure prediction
RoseTTAFold [40] - SE(3)-Transformer [41] 130M 2.03 GB
DP
GA
protein structure prediction
AlphaFold 3 [42] - Pairformer [42] Unreported - Unreported
biomolecular complex
structure prediction
OpenFold [43] - Evoformer 93M 1.45 GB
DP (ZeRO-2)
mixed-precision
offloading
GC, GA
protein structure prediction
FastFold [44] - Evoformer 93M 1.45 GB
DP
DAP [44]
protein structure prediction
ScaleFold [45] - Evoformer 97M 1.52 GB
DP
DAP
protein structure prediction
ESMFold [46] ESM-2 15B Transformer [17] 15B 240 GB DP (FSDP [47]) protein structure prediction
xTrimoPGLM [48] xTrimoPGLM-100B Transformer 100B 1.56 TB
DP (ZeRO-1)
PP (1F1B)
TP (Megatron-LM)
mixed-precision
GC
protein understanding
protein generation
ESM3 [49] - Transformer 1.4B / 7B / 98B 1.53 TB
DP (FSDP)
mixed-precision
protein reasoning
protein generation
Medicine BioGPT [50] GPT-2 Medium [25] Transformer 347M 5.42 GB
DP
GA
biomedical text understanding
biomedical text generation
Med-PaLM [3]
PaLM 540B [28]
Flan-PaLM 540B [26]
Transformer 540B 8.44 TB
DP (ZeRO-3)
TP
GC
medical question answering
medical reasoning
Med-PaLM 2 [4] PaLM 2 340B [29] Transformer 340B 5.31 TB Unreported
medical question answering
medical reasoning
Med-PaLM M [5] PaLM-E [51] Transformer 12B / 84B / 562B 8.78 TB
DP (ZeRO-3)
TP
GC
biomedical generalist
BiomedGPT [52] OFA [53] Transformer 33M / 93M / 182M 2.84 GB
DP (PyTorch DDP)
mixed-precision
biomedical generalist
Meditron [39] Llama-2 [19] Transformer 7B / 70B 1.09 TB
DP, PP
TP (Megatron-LM)
medical reasoning
HuatuoGPT [54]
Baichuan-7B
Ziya-LLaMA-13B [55]
Transformer 7B / 13B 208 GB DP (ZeRO) medical consultation
HuatuoGPT-II [56]
Baichuan2 [57]
Yi-34B [58]
Transformer 7B / 13B / 34B 544 GB DP (ZeRO) medical consultation
@@
PharmGPT [60] Llama-2 Transformer 3B / 13B / 70B 1.09 TB
DP+PP+TP
text understanding
text generation
Chemistry ChemBERT [61] BERT-Base Transformer 110M 1.72 GB Unreported
product extraction
reaction role labeling
@@
Chemformer [9] BART [63] Transformer 45M / 230M 3.59 GB
DP (ZeRO-2)
mixed-precision
reaction prediction
molecular optimization
molecular property prediction
MolGen [7] BART Transformer 355M 5.55 GB
DP (ZeRO-2)
mixed-precision
GA
molecule generation
@@
FuXi [12] - Swin transformer [66] 4.5B 72 GB
DP (FSDP)
mixed-precision
GC
weather forecast
ClimaX [13] - ViT [67] Unreported -
DP
mixed-precision
weather forecast
climate projection
Aurora [14] - Swin transformer 1.3B 20.80 GB
DP
mixed-precision
GC
atmospheric prediction
Geoscience K2 [68] Llama-7B [19] Transformer 7B 112 GB
DP (ZeRO-3)
mixed-precision
GA
foundational model
GeoGalactica [69] Galactica-30B [70] Transformer 30B 480 GB DP+PP+TP+SP foundational model
JiuZhou [71]
Mistral-7B [72]
Qwen1.5-14B [37]
Transformer 7B / 14B 224 GB
DP (ZeRO-2)
mixed-precision
GA
foundational model
  • DP: data parallelism,   TP: tensor parallelism,   PP: pipeline parallelism,   SP: sequence parallelism,   GA: gradient accumulation,   GC: gradient checkpointing
2.2.1 Biology
@@
Protein structure prediction represents a significant challenges in biochemistry, as high-precision predictions are essential for drug discovery and protein design. AlphaFold 2 [2] is a DL-based tool for predicting protein structures. It can predict the three-dimensional structure of proteins based on their amino acid sequences.
@@
The architecture of AlphaFold 2 is built upon variants of Transformer. Its main structure features an input module that receives the amino acid sequence of the target protein and searches for homologs within a sequence database to perform multiple sequence alignment (MSA). The Evoformer module then processes the MSA representation and pair representation, facilitating the exchange and updating of information through Transformer layers. Finally, the structural module converts the abstract representation of proteins into 3D atomic coordinates [73]. AlphaFold 2 has significantly influenced the advancement of biotechnology. However, its training process and datasets are not fully open source. Furthermore, despite the AlphaFold 2 model having a relatively small parameter size (about 93M), the memory consumed by intermediate activations is substantial due to its architecture, and peak memory usage increases cubically with the length of the input sequence.
@@
Although AlphaFold2 and RoseTTAFold [40] achieved breakthroughs in atomic resolution structure prediction, they rely on MSA and similar protein structure templates to achieve better performance. Based on the research of OpenFold [43], Lin et al. [46] further explored the ability of language model and proposed ESMFold. It utilizes the internal representation of language model to replace the expensive network module for processing MSA with a transformer module for processing sequences. This simplification greatly improves the speed of ESMFold, which is much higher than MSA-based models.
@@
The protein language model (PLM) has achieved remarkable success in learning biological information from protein sequences. Most existing models are constrained by autoencoding or autoregressive pre-training objectives, making it challenging to address both protein understanding and generation tasks. Chen et al. [48] explored the possibility of jointly optimizing these two training objectives through an innovative framework and trained a 100B PLM (xTrimoPGLM). Inspired by ESMFold, they combined the folding module with xTrimoPGLM and developed xT-Fold, a high-performance 3D structure prediction tool. The performance of xT-Fold is superior to other PLM-based models on benchmarks such as CAMEO and CASP15.
@@
In May 2024, AlphaFold 3 was released [42]. The overall architecture of Alphafold 3 is similar to AlphaFold 2, yet it demonstrates enhanced prediction accuracy and efficiency. dditionally, the application scope of Alphafold 3 has been expanded. It can predict the structure of a single protein, the complex structure of protein complexes, proteins, nucleic acids, small molecules, etc.
2.2.2 Medicine